For the next two questions, consider three acid solutions represented below. Acid I Acid II Acid III 31, which acid is the strongest? A. acid I B. acid I C. acid III D. the acids are similar in strength -32, which solution has the highest pH? A. acid I B. acid II C. acid II D. the solutions have similar pH ー33. which of the following salts will form a basic aqueous solution? A. NH Br B. Ca(NO)2 C. CSF D. AICI, 34. In the relationship Δ G-_nFE, what is the value of n for the reaction shown below? 3 Cu(aq) 2 Al(s) 3 Cu(s) + 2 A (aa) A. 1 B. 2 C. 3 D. 6 35. The ideal vant Hoff factors for CH3CO2Na, K2SO4 and (NH4)s(PO4) are, respectively... A. 2, 3 and 4 B. 3, 2 and 4. C. 2, 3 and 5. D. 3, 3 and 4.
Answer #1

Q31.

stronger acids will donate 100% therefore, expect more H3O+ in solution

weak acids will not donate 100% protons, therefore, expect some H3O+ and H2O balance

From the list,

choose Acid II, since there is only A- and H3O+ which implies there is 100% dissociation

Q32.

higher pH is that which is lower in H+, since

pH = -log(H+)

therefore, the more H+ in solution, the lower pH value

frm the list, the weaker acid

this must be Acid I, since it has the least H3O+ molecules per HA (ratio)

then, Acid I has higher pH

Q33.

basic solution = must have OH- in solution higher than H+

from the list:

NH4Br = NH4+ + Br-, this is acidic since NH4+ = NH3 + H+

Ca(NO3)2 = Ca2+ + 2NO3- slightly acidic / neutrla

CsF = Cs+ + F-

F- + H2O = HF + OH-

note that OH- is in solution, therefore this is BASIC

chose CsF

Q34.

n = number of electrons being transferred

the electron balance:

Copper goes from +2 to 0; loses 2 electron pe mol of Cu

1 mol of Cu = 2 electrons

3 mol of Cu = 2x3 = 6 mol of e-

then

n = 6

Q35.

this value is the total amount of ions (count)

so

CH3CO2Na = CH3CO2- + Na ( 2ions)

K2SO4 = 2K+ + SO4-2 (3 ions)

(NH4)3(PO4) = 3NH4+ and 1 PO4- = 3+1 = 4

answer is 2,3,4

choose A